How many joules are required to raise the temperature of 450 mL of water from 30° C to 100 °C?



1
Expert's answer
2022-03-29T11:26:04-0400

Solution:

The density of water is around approximately 1 g/cm3 or 1 g/mL

Therefore,

Mass of water = Density of water × Volume of water

Mass of water = (1 g mL−1) × (450 mL) = 450 g


q = m × C × ΔT

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C−1 g−1)

ΔT = change in temperature (°C)


m = 450 g

C = 4.184 J °C−1 g−1 (for water)

ΔT = 100°C − 30°C = 70°C


Thus:

q = (450 g) × (4.184 J °C−1 g−1) × (70°C) = 131796 J

q = 131796 J


Answer: 131796 joules are required